Band A charges $600 to play for the entire evening.
Band B charges $350 plus $1.25 for each ticket.
I have two questions:
1. How would I write an equation for each band?
2. How do I find the number of tickets for which the cost of the two bands will be equal?

Answer by oberobic(2304)

Band A charges $600.
Total Cost = 600
.
Band B charges 350 + $1.25 per person
Total Cost = 350 + 1.25y
, where y = number of tickets sold.
.
To find when they're equal, simply set Total Cost = Total Cost, which we can model with:
600 = 350 + 1.25y
250 = 1.25y
1.25y = 250
y = 200
.
So their costs are equal when there are exactly 200 people in attendance.
.
Remember to check your work!
.
350 + 1.25*200 = 350+250 = 600